Hakara Company has been using direct labor costs as the basis for assigning overhead to its many products. Under this allocation system, product A has been assigned overhead of $28.26 per unit, while product B has been assigned $11.09 per unit. Management feels that an ABC system will provide a more accurate allocation of the overhead costs and has collected the following cost pool and cost driver information:

Cost Pools Activity Costs Cost Drivers Activity Driver Consumption
Machine setup $ 128,000 Setup hours 2,000
Materials handling 84,000 Pounds of materials 14,000
Electric power 33,000 Kilowatt-hours 33,000

The following cost information pertains to the production of A and B, just two of Hakara's many products:

A B
Number of units produced 4,000 10,000
Direct materials cost $ 36,000 $ 25,000
Direct labor cost $ 34,000 $ 36,000
Number of setup hours 100 200
Pounds of materials used 1,000 1,000
Kilowatt-hours 2,000 4,000

Required:

1. Use activity-based costing to determine a unit cost for each product. (Round your final answers to 2 decimal places.)

Cost per Unit for:

Product A.

Product B.
Answer #1

Answer

  • Requirement 1 with working

Product A

Product B

Direct Material Cost

$36,000

$25,000

Direct Labor Cost

$34,000

$36,000

Machine setup cost

$6,400

$12,800

Material handling cost

$6,000

$6,000

Electric power cost

$2,000

$4,000

Total Cost

$84,400

$83,800

Total Units

4000

10000

Unit Cost

$21.10

$8.38

--Working

Product A

Product B

Direct Material Cost

36000

25000

Direct Labor Cost

34000

36000

Machine setup cost

=(128000/2000)*100

=(128000/2000)*200

Material handling cost

=(84000/14000)*1000

=(84000/14000)*1000

Electric power cost

=(33000/33000)*2000

=(33000/33000)*4000

Total Cost

$84,400

$83,800

Total Units

4000

10000

Unit Cost

$21.10 ($84400 / 4000)

$8.38 ($83800 / 1000)